From the equation 3c2h2 gives c6h6. Find the volume of acetylene for the manufacture of 3 mol of benzene
 

Grade:12

1 Answers

Arun
4 years ago
Dear student
 
for 1 mole of Benzene = 3* 22.4 litre is used 
 
hence for 3 moles of benzene = 3 * 3*  22.4 = 201.6 litres
 
Hope it helps